OpenStudy (anonymous):

7/9 with denominator of 27 How do I write this out to solve it?

OpenStudy (anonymous):

do you have any equation?

OpenStudy (anonymous):

nope thats the question itself.

OpenStudy (anonymous):

7/9/27

OpenStudy (lgbasallote):

hint: \[\huge \frac{\frac 79}{27} \implies \frac{\frac 79}{\frac{27}1} \implies \frac 79 \div \frac {27}1 \implies \frac 79 \times \frac 1{27}\] does that help?

OpenStudy (anonymous):

Is this the reciprocal method?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

243 is the answer

OpenStudy (anonymous):

i guess but do the math

OpenStudy (lgbasallote):

you should have a fraction as an answer o.O

OpenStudy (anonymous):

That question ended up changing to 5/7 with denominator of 42 and as an answer I got 30/42 following your example.

OpenStudy (anonymous):

I mean I didn't get 30/42. I got 5/294

OpenStudy (anonymous):

0.02880658436 this is the answer

OpenStudy (anonymous):

7/243 = 0.02880658436

OpenStudy (anonymous):

The correct answer is 30/42

OpenStudy (lgbasallote):

i think your question is unclear then

OpenStudy (anonymous):

then you wrote the question wrong

OpenStudy (anonymous):

Then I have no idea.

OpenStudy (lgbasallote):

is the question \[\huge \frac{\frac 79}{27}\]

OpenStudy (anonymous):

No it got changed.

OpenStudy (anonymous):

and I already got it wrong on my online homework so i cant go back and change the answer.

OpenStudy (anonymous):

But i'd like to have more example problems.